OpenStudy (anonymous):

A coin is tossed and a spinner with four equal sections numbered 1 through 4 is spun. What is the probability that the coin will land with heads up and the spinner will stop on 3? A. 1/8 B. 1/6 C. 1/3 D. 3/4

OpenStudy (xguardians):

1/2 x 1/4 A is answer.

OpenStudy (anonymous):

P(getting a Head) * P(Spinning a 3)=1/2 * 1/4= 1/8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime ) @CUTIEGIRL01